OpenStudy (anonymous):

Given the equation y − 3 = (x + 6) in point-slope form, identify the equation of the same line in slope-intercept form. x − 2y = −12 y = x y = x + 6 y = x + 9

OpenStudy (texaschic101):

slope intercept form is y = mx + b y - 3 = x + 6 (add 3 to both sides) y - 3 + 3 = x + 6 + 3 (combine like terms) y = x + 6 + 3 y = x + 9 < -- slope intercept form